OpenStudy (anonymous):

5/(x+4)-6/(x-3)=4/4 Find the values of x that could result in an extraneous solution

OpenStudy (owlcoffee):

so we have: \[\frac{ 5 }{ x+4} - \frac{ 6 }{ x-3 }= \frac{ 4 }{ 4 }\] Now in order to find a extraneous solution of this, we must find a value of x that makes both denominators equal 0.